OpenStudy (autumnbk98):

Solve for x: 2 over 5 plus 3 over 5 x equals quantity x plus 5 over 10.

OpenStudy (sshayer):

\[\frac{ 2 }{ 5 }+\frac{ 3 }{ 5x }=\frac{ x+5 }{ 10 }\] \[\frac{ 3 }{ 5x }=\frac{ x+5 }{ 10 }-\frac{ 3 }{ 5 }=\frac{ x+5-6 }{ 10 }=\frac{ x-1 }{ 10 }\] cross multiply and solve for x,either by quadratic formula or by factoring.

OpenStudy (sshayer):

correction write 2/5 instead of 3/5.

OpenStudy (autumnbk98):

can you help me with the factoring?

OpenStudy (sshayer):

\[\frac{ 3 }{ 5 x }=\frac{ x+5 }{ 10 }-\frac{ 2 }{ 5 }=\frac{ x+5-4 }{ 10 }=\frac{ x+1 }{ 10 }\] \[5x \left( x+1 \right)=3 \times 10\] \[5 x^2+5 x=30\] \[5 x^2+5 x-30=0\] divide by 5 what do you get?

OpenStudy (autumnbk98):

-5

OpenStudy (autumnbk98):

i mean -3

OpenStudy (sshayer):

after dividing by 5 \[x^2+x-6=0\]

OpenStudy (sshayer):

\[x^2+(3-2)x-6=0\] \[x^2+3x-2x-6=0\] now take common and make factors.

OpenStudy (autumnbk98):

what do you mean by take common?